The following measurements were recorded for the drying time, in hours, of a certain brand of latex paint:


3.4 2.5 4.8 2.9 3.6

2.8 3.3 5.6 3.7 2.8

4.4 4.0 5.2 3.0 4.8


Assuming that the measurements represent a random sample from a normal population, find a 95% prediction interval for the drying time for the next trial of the paint.


1
Expert's answer
2021-11-01T08:47:41-0400

Lets mark u' as prediction interval

"u' = (u-Cr*s, u+Cr*s)" , where u - sample mean, Cr - testing critical value, s - sample standart deviation

"u = {\frac {3.4+...+4.8} {15}}={\frac {56.8} {15}}=3.79"

"s^2={\frac {(3.4-3.79)^{2}+...+(4.8-3.79)^{2}} {14}}=0.986\to s=0.993"

Since the sample population has normal distribution, it is appropriate to calculate critical value using t-statistic with 15 - 1 = 14 degrees of freedom

"P(t(14) > Cr) = {\frac a 2}=0.025" , where a - level of significance. We use "{\frac a 2}" cause we need a two-sided test

Now the value of the Cr can be found from tables of t-statistic

Сr = 2.145

So, "u-Cr*s=3.79-2.145*0.93=1.66" , "u+Cr*s=5.92"

The 95% prediction interval is (1.66, 5.92)